# Attention as an Adaptive Filter

### v1

@@ 5.6 Dynamical Systems Perspectives on Attention
Stochastic Recurrent Networks (STORNs) combine RNNs with time-dependent latent variables and variational inference to model multimodal temporal dynamics (Bayer and Osendorfer, 2015). Neural ODEs treat hidden states as continuous-time dynamics parameterized by a neural network and solved with an ODE solver (Chen et al., 2019). Neural Processes (NPs) and Neural ODE Processes (NDPs) extend this framework to learn distributions over functions or ODE trajectories, enabling uncertainty-aware predictions (Garnelo et al., 2018; Norcliffe et al., 2021). Rubanova et al. introduce ODE-RNNs and Latent ODEs, continuous-time models that handle irregularly-sampled time series without discretization or imputation (Rubanova et al., 2019). Physics-informed neural networks (PINNs) integrate differential equation constraints into neural networks to enforce physically consistent solutions (Raissi et al., 2017).
@@
Attentive Neural Processes (ANPs) introduce attention over context sets, while Attentive Differential Networks (ADNs) handle irregularly sampled data using continuous-time attention (Kim et al., 2019; Chien and Chen, 2023). ACE-NODE and ANCDE integrate attention into ODE or controlled differential equation frameworks, learning attention dynamics continuously rather than via standard query-key comparisons (Jhin et al., 2021a, b). Neural ODE Transformers and Hopfield Networks reinterpret attention as evolving or energy-based dynamical systems (Tong et al., 2025; Ramsauer et al., 2021). Continuous-discrete state space models and Structured Inference Networks combine latent continuous dynamics with discrete observations and approximate Bayesian filtering (Ansari et al., 2023; Krishnan et al., 2016).
Geshkovski et al. model Transformers as interacting particle systems evolving on the unit sphere, and find that self-attention induces nonlinear dynamics leading to clustering of token embeddings (Zhang et al., 2021; Geshkovski et al., 2024). Molina interprets Transformers geometrically, showing that LayerNorm constrains hidden states to a hypersphere and attention moves “word particles” along this surface (Molina, 2023). Brody et al. show that LayerNorm shapes Transformer hidden states geometrically—projecting them onto a hyperplane and scaling them—which enhances the expressiveness of attention (Brody et al., 2023).
5.7 Diffusion, Score-Based Models, and Neural SDEs
@@
Song et al. showed that diffusion models can be formulated as continuous-time stochastic differential equations (SDEs), where learning the time-dependent score allows high-quality sampling via SDE solvers (Song et al., 2021). FIM-SDE estimates drift and diffusion functions from noisy, sparse SDE observations by learning a generalizable mapping from trajectories to underlying dynamical system parameters (Seifner et al., 2025). Tzen and Raginsky show that deep latent Gaussian models converge in the infinite-layer limit to Itô SDEs with neural-network-parameterized drift and diffusion (Tzen and Raginsky, 2019). Wang and Yao combine neural processes with neural SDEs to improve uncertainty estimation on noisy or incomplete data, using attention-based NPs for multidimensional inputs (Wang and Yao, 2021).
5.8 Structuring Attention with Priors and Positional Biases
The original Transformer lacks built-in modeling of relative or absolute positions, relying instead on added absolute positional encodings (Vaswani et al., 2023).
Several methods have proposed extending self-attention to efficiently incorporate relative position information. For example, Shaw et al. introduce a mechanism to encode relative distances between tokens directly into the attention computation (Shaw et al., 2018). Transformer-XL introduces relative positional encodings, enabling the model to generalize to sequences longer than those seen during training (Dai et al., 2019). Su et al. apply a position-dependent rotation to each dimension of the query and key vectors before computing their dot products. This rotation encodes the relative distance between positions directly into the attention mechanism (Su et al., 2023).
A number of works reinterpret sequence modeling as a controlled, time-aware aggregation of past information, where learned or adaptive decay mechanisms regulate the influence of prior inputs over time. These models typically rely on heuristic, learned gating or decay functions for memory control, whereas we derive update weights from robust, data-driven precision estimates grounded in Bayesian filtering.
The Retentive Network (RetNet) introduces a retention mechanism that blends each token’s hidden state with an exponentially decaying summary of previous tokens. RetNet uses multiple attention heads, each having its own learned exponential decay rate, enabling multi-scale memory of different temporal lengths. Attention is applied with decay masks that enforce causality and decay over relative positions, while a gating mechanism modulates the retained summaries to improve expressivity and control the contribution of past information (Sun et al., 2023).
GateLoop generalizes linear RNNs with data-controlled, time-varying state transition gates and introduces a surrogate attention mechanism where each input contributes via learned, input-dependent decay factors, capturing content and relative position (Katsch, 2024). Gated Linear Attention (GLA) uses data-dependent, time-varying gates to modulate hidden state updates, combining fast, recurrent linear attention with input-adaptive, relative position-aware decay (Yang et al., 2024). The Real-Gated Linear Recurrent Unit (RG-LRU) introduces input and recurrence gates into a stable linear recurrent update, avoiding complexity from softmax attention or continuous-time dynamics. Griffin interleaves RG-LRU blocks with lightweight local attention (De et al., 2024). The Dynamically Context-Sensitive Time-Decay Attention mechanism learns a time-decay function to adjust attention weights (Su et al., 2018). This approach focuses on temporal proximity rather than integrating known system dynamics. Performer approximates softmax attention with positive-definite kernel methods to achieve scalable, linear-time self-attention for long sequences (Choromanski et al., 2022).
Several papers have explored element-wise (in the embedding dimension) attention mechanisms. McCarter investigates replacing scaled dot-product attention with negative-log Euclidean distance (McCarter, 2023). Elliptical Attention replaces the traditional dot-product attention with a Mahalanobis distance-based mechanism (Nielsen et al., 2024). Feng uses an exponentiated Euclidean distance as a similarity metric (Feng, 2025). Bacochina et al. replace matrix multiplications with learned spatial weight arrays applied directly via element-wise multiplication, focusing attention along rows and columns (Bacochina and de Souza, 2023). Wu et al. introduce the DA-Transformer, which reweights attention scores using learnable functions of token distances (Wu et al., 2021)
@@ 5.9 Modeling Uncertainty in Neural Networks
Other approaches explicitly model uncertainty in neural networks and attention mechanisms. Foundational works such as (MacKay, 1992) and (Neal, 1996) laid the groundwork for modern developments in approximate Bayesian inference in deep learning.
Chung et al. introduce the Variational Recurrent Neural Network (VRNN), which incorporates latent random variables into RNN hidden states to represent and propagate uncertainty in sequential data (Chung et al., 2016). Fraccaro et al. introduce the stochastic recurrent neural network, a sequential model combining deterministic RNN states with stochastic latent variables, using structured variational inference to propagate uncertainty over time and improve sequence modeling (Fraccaro et al., 2016).
@@ 5.10 Structured Low-Rank Covariances for Efficient Propagation
5.11 Unrolling Optimization Algorithms and Connections to Attention
Unrolling iterative optimization algorithms into trainable networks began with frameworks like LISTA, which expresses the ISTA sparse-coding algorithm as a neural network with learned parameters (Gregor and LeCun, 2010). Deep Equilibrium Models (DEQs) extend this idea by solving a single equilibrium layer via root-finding, effectively modeling infinite-depth networks with constant memory (Bai et al., 2019). Wang et al. study how attention can also be viewed as an unrolled iterative process: multi-head self-attention corresponds to iterative subspace denoising, compressing inputs into a low-rank manifold (Wang et al., 2025).
@@ 6 Discussion and Conclusion
This work explores a connection between classical control theory and attention mechanisms by framing attention through the lens of linear dynamical systems and adaptive filtering. We show that attention arises naturally as the solution to a maximum likelihood estimation problem under a linear stochastic differential equation with structured system matrices and noise covariances, and discrete observations. In this view, the attention update implements a parallelized form of adaptive Kalman Filtering: each key-value pair is interpreted as a noisy observation of a latent state, and a precision-weighted Bayesian update aggregates evidence across all observations to refine the state estimate in a parallel manner. To the best of our knowledge, this view of attention as the MLE for a structured linear SDE has not appeared in the literature on attention mechanisms.
Several promising directions remain for future work. It should be straightforward to extend the method to the LQG controller, and the framework could also be adapted to reinforcement learning settings, where the dynamics-aware attention mechanism might provide more sample-efficient policy evaluation and state estimation in partially observed or noisy environments.
@@
In this work, we assume the output matrix $\mathbf{C}$ is square and invertible, so the latent state is fully observable. We can account for partially observed states with projections or pseudo-inverses to map between measurement and latent spaces.
The layer-wise updates in our model perform adaptive denoising, which parallels the iterative refinement in diffusion models. Future work could investigate these connections. Finally, it would be worthwhile to explore whether linking attention to system dynamics and measurement confidence can improve model interpretability.
@@ Appendix A Background
A.1 Linear Systems
A continuous-time linear system is defined, in state-space form, as:
$\mathbf{\dot{x}}(t)=\mathbf{A}(t)\mathbf{x}(t)+\mathbf{B}(t)\mathbf{u}(t)$
$\mathbf{y}(t)=\mathbf{C}(t)\mathbf{x}(t)+\mathbf{D}(t)\mathbf{u}(t)$
The signals:
$\textbf{x}(t):[0,\infty)\to\mathbb{R}^{N},\quad\textbf{u}(t):[0,\infty)\to\mathbb{R}^{M},\quad\textbf{y}(t):[0,\infty)\to\mathbb{R}^{P}$
are known as the state, input, and output, respectively. In control applications, $\mathbf{u}(t)$ is usually a feedback control term. $\mathbf{A}(t),\mathbf{B}(t),\mathbf{C}(t),\mathbf{D}(t)$ are known as the state matrix, input matrix, output matrix, and direct transition matrix, respectively.
Letting $\textbf{x}(t_{0})=\textbf{x}_{0}$, the unique solution to the above system is given by:
$\textbf{x}(t)=\mathbf{\Phi}(t,t_{0})\textbf{x}_{0}+\int_{t_{0}}^{t}\mathbf{\Phi}(t,\tau)\mathbf{B}(\tau)\textbf{u}(\tau)\,d\tau,$
$\textbf{y}(t)=\mathbf{C}(t)\mathbf{\Phi}(t,t_{0})\textbf{x}_{0}+\int_{t_{0}}^{t}\mathbf{C}(t)\mathbf{\Phi}(t,\tau)\mathbf{B}(\tau)\textbf{u}(\tau)\,d\tau+\mathbf{D}(t)\textbf{u}(t),$
where $\mathbf{\Phi}(t,t_{0})$, the state transition matrix, is the unique solution to
$\frac{d}{dt}\mathbf{\Phi}(t,t_{0})=\mathbf{A}(t)\mathbf{\Phi}(t,t_{0}),$
@@ A.1 Linear Systems
A linear time-invariant (LTI) system is one in which the matrices $\mathbf{A},\mathbf{B},\mathbf{C},\textbf{D}$ are fixed, real-valued matrices, (i.e. not functions of time): $\mathbf{A}\in\mathbb{R}^{N\times N}$, $\mathbf{B}\in\mathbb{R}^{N\times M}$, $\mathbf{C}\in\mathbb{R}^{P\times N}$, $\mathbf{D}\in\mathbb{R}^{P\times M}$.
An autonomous LTI system $\mathbf{\dot{x}}(t)=\mathbf{A}\mathbf{x}(t)$ is asymptotically stable (i.e. $\mathbf{x}(t)$ approaches a stable equilibrium point as $t\rightarrow\infty$) if $\mathbf{A}$ is Hurwitz, that is if every eigenvalue of $\mathbf{A}$ has strictly negative real part.
In an LTI system, the state transition matrix is given by:
$\mathbf{\Phi}(t,t_{0})=e^{\mathbf{A}(t-t_{0})},$
where the matrix exponential $e^{\mathbf{M}}$ is defined as:
$e^{\mathbf{M}}=\sum_{k=0}^{\infty}\frac{1}{k!}\mathbf{M}^{k}.$
If $\mathbf{A}$ is diagonalizable, i.e. $\mathbf{A}=\mathbf{S}\bm{\Lambda}\mathbf{S}^{-1}$ for some invertible $\mathbf{S}$ and diagonal $\bm{\Lambda}$, then the matrix exponential can be computed cheaply using: $e^{\mathbf{A}}=\mathbf{S}e^{\bm{\Lambda}}\mathbf{S}^{-1}$. See also, standard references such as (Antsaklis, 1997), (Skogestad and Postlethwaite, 2005).
A.2 Stochastic Processes
Many real-world systems are subject to uncertainty, which can be modeled as stochastic disturbances or process noise. A continuous-time linear time-invariant system with additive Gaussian noise is described by the Itô stochastic differential equation (SDE):
$d\mathbf{x}(t)=\mathbf{A}\mathbf{x}(t)\,dt+\mathbf{B}\,d\mathbf{w}(t),\qquad\mathbf{z}(t_{k})=\mathbf{C}\mathbf{x}(t_{k})+\mathbf{v}(t_{k}),$

### v2

@@ 6.8 Dynamical Systems Perspectives on Attention
Stochastic Recurrent Networks (STORNs) combine RNNs with time-dependent latent variables and variational inference to model multimodal temporal dynamics (Bayer and Osendorfer, 2015). Neural ODEs treat hidden states as continuous-time dynamics parameterized by a neural network and solved with an ODE solver (Chen et al., 2019). Neural Processes (NPs) (Garnelo et al., 2018) and Neural ODE Processes (NDPs) (Norcliffe et al., 2021) extend this framework to learn distributions over functions or ODE trajectories, enabling uncertainty-aware predictions. Rubanova et al. introduce ODE-RNNs and Latent ODEs, continuous-time models that handle irregularly-sampled time series without discretization or imputation (Rubanova et al., 2019).
@@
Chien and Chen (2021) introduce continuous-time attention, integrating attention mechanisms with neural differential equations to operate over latent trajectories rather than discrete hidden states, enabling effective modeling of irregularly-sampled sequences (Chien and Chen, 2021). Attentive Neural Processes (ANPs) extend Neural Processes by incorporating attention to allow each input to selectively attend to relevant context points (Kim et al., 2019). Chen at al. introduce self-modulating attention , which extends classical attention to continuous-time sequences by adaptively reweighting based on both sequential order and temporal distance (Chen et al., 2022). ACE-NODE and ANCDE integrate attention into ODE or controlled differential equation frameworks, learning attention dynamics continuously rather than via standard query-key comparisons (Jhin et al., 2021a, b). Neural ODE Transformers and Hopfield Networks reinterpret attention as evolving or energy-based dynamical systems (Tong et al., 2025; Ramsauer et al., 2021). Continuous-discrete state space models and Structured Inference Networks combine latent continuous dynamics with discrete observations and approximate Bayesian filtering (Ansari et al., 2023; Krishnan et al., 2016).
@@
Wang and Yao combine neural processes with neural SDEs to improve uncertainty estimation on noisy or incomplete data, using attention-based NPs for multidimensional inputs (Wang and Yao, 2021). FIM-SDE estimates drift and diffusion functions from noisy, sparse SDE observations by learning a generalizable mapping from trajectories to underlying dynamical system parameters (Seifner et al., 2025). Shen and Cheng propose a Neural SDE framework that models continuous-time sequences with neural network–parameterized drift and diffusion, using a decoupled maximum likelihood scheme for efficient, robust training (Shen and Cheng, 2025).
@@ 6.9 Modeling Uncertainty in Neural Networks
Other approaches explicitly model uncertainty in neural networks and attention mechanisms. Chung et al. introduce the Variational Recurrent Neural Network (VRNN), which incorporates latent random variables into RNN hidden states to represent and propagate uncertainty in sequential data (Chung et al., 2016). Fraccaro et al. introduce the stochastic recurrent neural network, a sequential model combining deterministic RNN states with stochastic latent variables, using structured variational inference to propagate uncertainty over time and improve sequence modeling (Fraccaro et al., 2016).
@@ 6.10 Structured Low-Rank Covariances for Efficient Propagation
6.11 Neural Network Layers as Unrolled Iterative Algorithms
Unrolling iterative optimization algorithms into trainable networks began with frameworks like LISTA, which expresses the ISTA sparse-coding algorithm as a neural network with learned parameters (Gregor and LeCun, 2010). Deep Equilibrium Models (DEQs) extend this idea by solving a single equilibrium layer via root-finding, effectively modeling infinite-depth networks with constant memory (Bai et al., 2019). Wang et al. study how attention can also be viewed as an unrolled iterative process: multi-head self-attention corresponds to iterative subspace denoising, compressing inputs into a low-rank manifold (Wang et al., 2025). This is analogous to our framing of attention as a step of iteratively-reweighted least squares.
6.12 Geometric Views of Transformers
Geshkovski et al. model Transformers as interacting particle systems evolving on the unit sphere, and find that self-attention induces nonlinear dynamics leading to clustering of token embeddings (Zhang et al., 2021; Geshkovski et al., 2024). Molina interprets Transformers geometrically, showing that LayerNorm constrains hidden states to a hypersphere and attention moves “word particles” along this surface (Molina, 2023). Brody et al. show that LayerNorm shapes Transformer hidden states geometrically — projecting them onto a hyperplane and scaling them — which enhances the expressiveness of attention (Brody et al., 2023). Loshichilov et al. introduce nGPT, a Transformer variant that normalizes all embeddings and network weights on the hypersphere, enabling cosine-similarity-based updates with learnable per-dimension step sizes, resulting in faster convergence and more stable training (Loshchilov et al., 2025). Gupta et al. observe that LayerNorm can be interpreted geometrically as projecting a vector orthogonally to the uniform vector, normalizing it, and scaling by the dimensionality, making it irreversible and largely redundant compared to RMSNorm, which achieves similar hidden representations more efficiently (Gupta et al., 2025).
@@ 7 Discussion and Conclusion
This work explores a connection between classical control theory and attention mechanisms by framing attention through the lens of linear dynamical systems and adaptive filtering. We show that attention arises naturally as the solution to a maximum likelihood estimation problem under a linear stochastic differential equation with structured system matrices and noise covariances, and discrete observations. Unlike approaches that treat attention as a static kernel or a matrix factorization, we view attention as a parallelized form of adaptive Kalman Filtering: each key-value pair is interpreted as a noisy observation of a latent state, and a precision-weighted Bayesian update aggregates evidence across all observations to refine the state estimate in a parallel manner. To the best of our knowledge, this view of attention as the MLE for a structured linear SDE has not appeared in the literature on attention mechanisms.
Imposing an isotropic-plus-skew-symmetric state matrix and isotropic noise, our model achieves the same asymptotic time and memory complexity as standard attention. This result in a form of complex-valued rotary positional encoding equivalent to that found in xPos and RetNet, along with a query-key precision kernel that quantifies an uncertainty prior under the learned dynamics model.
We also introduce a Radial–Tangential SDE model, whose isotropic-plus-skew-symmetric state matrix and anisotropic noise in the token frame define the most expressive SDE structure that remains analytically tractable. This formulation leads directly to a maximum likelihood estimator on the hypersphere, which provides a novel, principled foundation for understanding norm-preserving attention mechanisms, shedding new light on the architectural properties of the Transformer.
Several promising directions remain for future work. A natural extension is to allow the state matrix and noise to depend on the state. Although a fully time-varying dynamics model would break the linear time-invariant (LTI) structure and would no longer allow a closed-form solution to the Differential Lyapunov Equation (DLE), this issue can be avoided by fixing the dynamics within each context window. In this way, the parameters may adapt across windows while remaining constant within a window, enabling content adaptivity without sacrificing computational efficiency.
Future work should also investigate the relationship between the precision-weighted update and concepts from robust and adaptive control, such as $H_{\infty}$ filtering. This could lead to attention mechanisms with inherent guarantees of stability and robustness against dynamic model mismatch, which is critical for trustworthy AI applications.
@@
Finally, the duality between attention weights and the Kalman gain suggests a path toward dynamics-aware neural control policies, where the separation principle can be leveraged to achieve more sample-efficient policy evaluation and state estimation in noisy or partially observed reinforcement learning settings.
@@ Appendix A Background
A.1 Propagation of Uncertainty
Recall that the state at time $t_{j}$ can be written as:
$\mathbf{x}(t_{j})=e^{\mathbf{A}\Delta t_{ij}}\left(\mathbf{x}(t_{i})+\int_{0}^{\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{G}\,d\mathbf{w}(\tau)\right).$
The covariance of the integral term is:
$\mathbf{V}_{B}(\Delta t_{ij}):=\mathbb{E}\Bigg[\left(\int_{0}^{\Delta t_{ij}}e^{-\mathbf{A}\tau_{1}}\mathbf{G}\,d\mathbf{w}(\tau_{1})\right)\left(\int_{0}^{\Delta t_{ij}}e^{-\mathbf{A}\tau_{2}}\mathbf{G}\,d\mathbf{w}(\tau_{2})\right)^{\top}\Bigg]$
$\int_{0}^{\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{Q}e^{-\mathbf{A}^{\top}\tau}\,d\tau,\quad\mathbf{Q}=\mathbf{G}\mathbf{G}^{\top}$
where the forward-propagated state covariance is:
which is a solution of the differential Lyapunov equation:
with initial condition $\mathbf{V}_{F}(0)=\mathbf{0}$.
Recall that the estimate of the state is:
$\mathbf{\hat{x}}_{i}(t_{j}):=e^{\mathbf{-A}\Delta t_{ij}}\mathbf{x}(t_{j})=\mathbf{x}(t_{i})+\int_{0}^{\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{G}d\mathbf{w}(\tau).$
$\mathbf{x}(t_{i})=e^{\mathbf{A}(t_{i}-t_{j})}\mathbf{x}(t_{j})+\int_{t_{j}}^{t_{i}}e^{\mathbf{A}(t_{i}-\tau)}\mathbf{G}d\mathbf{w}(\tau)$
$e^{-\mathbf{A}\Delta t_{ij}}\left(\mathbf{x}(t_{j})+\int_{0}^{-\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{G}d\mathbf{w}(\tau)\right).$
$\mathbf{x}(t_{j})=e^{\mathbf{A}\Delta t_{ij}}\mathbf{x}(t_{i})-\int_{0}^{-\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{G}d\mathbf{w}(\tau)$
So:
We again form an estimate:
$\mathbf{\hat{x}}_{i}(t_{j})=e^{-\mathbf{A}\Delta t_{ij}}\mathbf{x}(t_{j})=\mathbf{x}(t_{i})-e^{-\mathbf{A}\Delta t_{ij}}\int_{0}^{-\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{G}d\mathbf{w}(\tau)$
@@ A.1 Propagation of Uncertainty
Recall that the measurement at time $t_{j}$ is:
$\mathbf{z}(t_{j})=\mathbf{C}e^{\mathbf{A}\Delta t_{ij}}\left(\mathbf{x}(t_{i})+\int_{0}^{\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{G}d\mathbf{w}(\tau)\right)+\mathbf{v}(t_{j}).$
The measurement process is distributed as:
$\mathbf{\hat{z}}_{i}(t_{j}):=\bm{\Phi}_{\mathbf{z}}(\Delta t_{ij})\mathbf{z}(t_{j})\quad\text{where}\quad\bm{\Phi}_{\mathbf{z}}(\Delta t):=\mathbf{C}e^{\mathbf{-A}\Delta t}\mathbf{C}^{-1}.$
In terms of the "pulled-forward" state estimate, this is:
$\mathbf{C}\mathbf{\hat{x}}_{i}(t_{j})+\bm{\Phi}_{\mathbf{z}}(\Delta t_{ij})\mathbf{v}(t_{j})$
Hence,
where the propagated measurement covariance is:
$\mathbf{V}^{C}(\Delta t_{ij})=\mathbf{C}\mathbf{V}(\Delta t_{ij})\mathbf{C}^{\top}+\bm{\Phi}_{\mathbf{z}}(\Delta t_{ij})\mathbf{R}\bm{\Phi}_{\mathbf{z}}^{\top}(\Delta t_{ij})$
A.2 Batch Maximum Likelihood Trajectory Estimation
Taking the derivative with respect to $\mathbf{x}_{i}$:
For the causal case:
$\mathbf{z}(t_{i+1})=\mathbf{C}e^{\mathbf{A}\Delta t}\mathbf{C}^{-1}\mathbf{z}^{C}(t_{i})+\mathbf{C}\int_{0}^{\Delta t}e^{-\mathbf{A}\tau}\mathbf{G}d\mathbf{w}(\tau)+\mathbf{v}(t_{i+1}),$
$\mathbf{z}^{+}(t_{i+1}):=\mathbb{E}[\mathbf{z}(t_{i+1})\mid\mathbf{\bar{z}}_{i}]=\mathbf{C}e^{\mathbf{A}\Delta t}\mathbf{C}^{-1}\mathbf{\bar{z}}_{i}.$
A.3 Closed-form Precision Matrix Calculation for Diagonalizable Matrices
Note that $(e^{-\bm{\Lambda}t}\mathbf{G}e^{-\bm{\Lambda}^{\dagger}t})_{kl}=\mathbf{G}_{kl}e^{-\lambda_{k}t}e^{-\lambda^{*}_{l}t}=\mathbf{G}_{kl}e^{-(\lambda_{k}+\lambda^{*}_{l})t}$. Hence, the $kl$th component of $\bm{\Lambda}_{\mathbf{V}B}(\Delta t)$ is:
$\bm{\Lambda}_{\mathbf{V}B}(\Delta t)_{kl}=\mathbf{G}_{kl}\int_{0}^{|\Delta t|}e^{-(\lambda_{k}+\lambda^{*}_{l})\tau}d\tau$
$\begin{cases}\mathbf{G}_{kl}\,\dfrac{1-e^{-(\lambda_{k}+\lambda^{*}_{l})|\Delta t|}}{\lambda_{k}+\lambda^{*}_{l}},&\lambda_{k}+\lambda^{*}_{l}\neq 0\\[5.69054pt] \mathbf{G}_{kl}\,|\Delta t|,&\lambda_{k}+\lambda^{*}_{l}=0\end{cases}$
For the causal case, the forward-propagated covariance is:
$\mathbf{S}\bm{\Lambda}_{\mathbf{V}}(|\Delta t_{ij}|)\mathbf{S}^{\dagger}$